OpenStudy (anonymous):

Integral[0,4] x/(sqrt 2x+1) dx

OpenStudy (lgbasallote):

\[\int_0^4 \frac{x}{\sqrt{2x+1}}dx?\]

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

interesting...

OpenStudy (lgbasallote):

are you allowed to use trig sub?

OpenStudy (anonymous):

i know that u= 2x+1

OpenStudy (lgbasallote):

nope

OpenStudy (anonymous):

idk about the top x

OpenStudy (lgbasallote):

that would make du = 2dx

OpenStudy (anonymous):

that what i thought

OpenStudy (lgbasallote):

have you discussed trigonemetric substitution?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

then use trig sub

OpenStudy (dumbcow):

no that works u = 2x+1 du = 2 dx then x = (u-1)/2 \[\rightarrow \frac{1}{4}\int\limits_{?}^{?}\frac{u-1}{\sqrt{u}} du\] then split that into 2 fractions

OpenStudy (lgbasallote):

hmm

OpenStudy (anonymous):

using the vdu= uv- int u dv

OpenStudy (dumbcow):

hmm no i wouldn't try integration by parts here

OpenStudy (anonymous):

@dumbcow how did u get the u-1?

OpenStudy (lgbasallote):

so it seems you can either do what the dumbcow did or trig sub

OpenStudy (lgbasallote):

\[2x + 1 = u\] \[u - 1 = 2x\] \[x = \frac{u-1}{2}\]

OpenStudy (anonymous):

let see both them lol

OpenStudy (anonymous):

@lgbasallote have you ever herped so much that you derped?

OpenStudy (lgbasallote):

herp derp?

OpenStudy (anonymous):

I don't think a trig sub would be useful here.......................................

OpenStudy (lgbasallote):

it would

OpenStudy (anonymous):

.................................at all

OpenStudy (lgbasallote):